# Pushing the Accuracy Limit of Foundation Neural Network Models with Quantum Monte Carlo Forces and Path Integrals

### v2

@@ V-B1 QMC Energy and Wavefunction Evaluation
We developed a highly optimized implementation of quantum Monte Carlo calculations in QMCPACK that leverages the massive parallelism of exascale hardware. Building on previous optimizations for material science applications, we specifically enhanced the Linear Combination of Atomic Orbitals (LCAO) branch to accelerate molecular applications on GPU architectures. Our implementation enables unprecedented calculation throughput, processing over 40,000 molecular configurations with QMC (20,251 with forces) on the Aurora supercomputer, achieving 124 PFLOPS sustained performance on 2048 nodes, and while using over 4500 walkers per rank (see Fig3).
@@
These implementations reduced kernel launch overhead by an order of magnitude and significantly improved memory throughput, particularly important for the LCAO approach used in molecular systems where the number of electrons is small. The architecture-independent optimizations for GPUs achieved 60% of theoretical peak performance for typical molecular configurations in our dataset. Additional batching strategies will allow us to reach
@@ V-B2 Single Determinant QMC Forces
For the Quantum Monte Carlo (QMC) layer of our quantum chemistry database generation protocol, we build upon the existing force evaluation formalism in QMCPACK, developed according to Filippi et al.’s  [25] approach. Rather than reimplementing the mathematical framework, we focused on optimizing the computational efficiency for exascale architectures by integrating the force evaluation with QMCPACK’s batched structure. This optimization enabled us to compute accurate interatomic forces for over 20,000 molecular configurations, achieving 62 PFLOPS sustained performance on Aurora with scaling efficiency of 30% up to 2048 nodes.
@@
By integrating force calculations with the batched structure of QMCPACK, we eliminated redundant computations and maximized hardware utilization, enabling force calculations at scales previously impossible with QMC methods. The performance improvements from our batched implementation make QMC force calculations practical for large-scale applications, opening new possibilities for highly accurate molecular dynamics simulations and foundation model training based on QMC-level accuracy.
@@ V-B3 Multideterminant QMC Forces
  • Utilizing the same flattened center indexing scheme used elsewhere in the code
@@
  • Minimize data transfers between CPU and GPU
  • Ensure optimal data locality for GPU processing
  • Reuse memory allocations across different parts of the calculation
@@
This implementation makes multideterminant QMC forces practically usable at scale for the first time. By completing, optimizing, and integrating this capability with our batched computation structure, we have created a pathway for generating higher-accuracy reference data for machine learning potentials, particularly for systems where electronic correlation effects are significant.
@@ V-B5 QMC Submission Engine on Aurora
QMCPACK calculations on Aurora were facilitated by a custom two-level job submission engine built on top of the Nexus[48] workflow system. To enable to maximize computational efficiency, molecular systems were batched together by electron count similarity as this supports greater execution uniformity among batched molecules. This top level parallelism, across batches, was complemented by further separation of molecules into sub-ensembles, each of which ran under a single top-level QMCPACK MPI process to maximize robustness of the simulations, providing a simple form of fault tolerance. The balance between these levels, as well as assignment of individual molecules to selected node resources, was optimized to simulaneously maximize job submission size, fault tolerance, and computational efficiency.
V-C Performance Analysis and Benchmarks
To demonstrate the computational efficiency and accuracy of our new Quantum Monte Carlo (QMC) implementation, we performed extensive performance analyses on several representative molecular systems. We primarily focused on water clusters (monomer, dimer, and trimer) due to their biological relevance and the well-known computational challenges they pose for accurate quantum chemical treatment. Additionally, we benchmarked a representative SPICE molecule with heavier atoms containing 16 electrons for force evaluation, as well as a larger, computationally demanding molecule with 176 electrons to examine scalability.
V-C1 Single-Determinant QMC Performance
@@
Figure 3 illustrates the substantial performance gains achieved with our new algorithm. We observe clear improvements across water clusters with electron counts ranging from 8 to 152 electrons. Quantitative speedups measured against the CPU baseline are summarized in Table I, highlighting consistent and significant GPU acceleration.
@@
V-C2 Resource Allocation Strategy
@@ V-C3 Multideterminant QMC Performance
We successfully executed multideterminant QMC calculations for the water monomer; however, performing analogous calculations on the water dimer exceeded our computational allocation limits, underscoring the extreme intensity of such computations. To achieve optimal performance with multideterminant wavefunctions, we allocated approximately 20 Aurora nodes per molecule and processed batches of 100 molecules per job (utilizing 2000 nodes per job submission). Our implementation reached a sustained computational throughput of 24 TFLOPS per node configuration on Aurora, corresponding to 25% of the peak theoretical performance.
These substantial accuracy gains justify the computational costs incurred, demonstrating the critical importance of multideterminant QMC methods in achieving chemically accurate reference datasets for training next-generation machine learning models.
V-D Large Scale Quantum Dynamics Simulations with a ”beyond DFT” Foundation Neural Network Model
@@
Using the complete basis-set QMC data produced in the project, we improved an early version of the FeNNix-Bio1 model [72] using transfer machine learning via the FeNNol library[69] to obtain a ”beyond DFT” improved approach. We then coupled the improved model to the GPU-accelerated massively parallel Tinker-HP molecular dynamics package [50, 1] and leverage simultaneously its Deep-HP (neural networks), Quantum-HP (quantum nuclear effects) modules while using enhanced sampling via non-supervised adaptive sampling molecular dynamics. We tested the new approach on both Jean-Zay (H100) and Aurora(Xe) machines on the full structure of the Satellite Tobacco Virus, a 1-million atom virus. The simulation encompassed 32-million particles (1-million atom X 32 beads) since it leverages Ring Polymer quantum molecular dynamics simulation. We extended the conformational space sampling with adaptive sampling replica to decompose the global exploration problem into a set of separate MD trajectories. These latter can be restarted thanks to a selective process in order to achieve sufficient phase-space sampling.
@@
    As a proof of concept, in order to improve further the quality of the DFT-based FeNNix-Bio1 model towards a ”beyond DFT” model, we leveraged transfer learning,[92] to include the high-quality, complete basis-set, DMC energies and forces present in the SPICE2(+)-ccECP dataset. In practice, starting from the model trained on the DFT part of the dataset, we reoptimize the energy heads (freezing the parameters of the embedding), on the QMC energies and forces with the same loss function as in the original training. As is common practice, we started with a smaller learning rate of $10^{-5}$ using the adabelief optimizer[107].
@@
    We coupled our newly developed JAX-based [16] FeNNoL library [69] that allows for the construction and inference of atomistic machine learning potentials such as FeNNix-Bio1[72] to the Tinker-HP molecular dynamics package[50, 1] via the Deep-HP interface[39]. By doing so, we leverage the Ring Polymer Molecular Dynamics (RPMD)[36] capabilities of Tinker-HP (Quantum-HP module [71]) in combination with FeNNix-Bio1 in order to recover quantum dynamics and statistics of the Satellite Tobacco Virus (STMV) that we previously studied [15] using polarizable molecular dynamics. Indeed RPMD approximates the Kubo-transformed time correlation function and can be used efficiently for determining accurate dynamical properties of condensed-phase molecular systems. It allows to use classical statistical physics to mimic quantum dynamics via fictitious ring polymer which consist in n copies (or beads) of the system particles connected by harmonic springs. The interested reader can refer to ref.[71] for details about RPMD and its implementation. Accounting for these has been shown to be mandatory to recover correct properties in that context and up to now FeNNix-BIo1 only resorted to more approximated methods such as adQTB[62] to mimic quantum dynamics. If adQTB is extremely fast, i.e. roughly similar computational cost as classical dynamics, and well suited for most of the FeNNix-BIo1 applications [72], the present RPMD implementation allows to reach an unprecedented accuracy on the FeNNiX-Bio1 quantum dynamics leading to an improved control on the chemical reactions, notably proton transfers. To ensure full accuracy at room temperature, we used 32 beads per atom in RPMD simulations. These latter leverage two levels of parallelism: one at the beads level and a second one within each bead. The choice of the STMV was motivated as it is a model system for studying viral assembly. Moreover, since it is includes a single-stranded RNA genome located within an icosahedral capsid, it is also a good model to study the structural dynamics stability of viruses[15]. Finally, to push further the sampling of the conformational space, we added on top of the discussed framework an additional layer of parallelism using automatized (unsupervised) adaptive sampling following described previously[40]. Here multiple iterations of independent molecular dynamics simulations replica are launched to ”mine” the conformational space. The framework is an iterative process and the initial conditions at each iteration are based on the results of previous iteration steps. The selection of the initial structures at each iteration follows an adaptive, fully automated, procedure enhancing the exploration of a low-dimensional space of slow variables. This unsupervised selection step has the advantage of suppressing the problem of the choice of the initial collective variable at the beginning of the simulation. In practice the procedure maximize the conformational space exploration by penalizing areas that have already been extensively visited. The choice of the number of replica is only limited by the number of available GPUs, making adaptive sampling particularly suited for exascale computing (see section of reference [40] for a detailed discussion. On Jean-Zay, each of the bead was parallelized on 8 H100 GPUs for a total of 8x32=256 H100s. Exploiting parallelism within each bead is mandatory to fit in the memory of each GPU thanks to memory distribution. All in all, we obtain 0.11 ns/day (mixed precision) per single adaptive sampling replica, making production of tens of nanoseconds per day of such quantum dynamics simulation accessible through adaptive sampling given enough resources. To our knowledge, achieving such level of quantum dynamics with a foundation model and grounded on extremely accurate quantum chemistry dataset on such a large biosystem is unprecedented. The Jean Zay machine having 1280 GPUs, we were able to launch 4 simultaneous replica leveraging 1024 GPUs, producing therefore 0.44 ns/day of adaptive quantum dynamics. Note that such adaptive setup can also be used sequentially to extend the number of replica that can be launched one after another enabling to accumulate simulation time. Overall, this computing scheme is therefore particularly suited for large exascale machine and further H100 optimization of the combined computing layers should possibly enable to reach the 1 ns/day threshold.
@@
    In perspective to run Tinker-HP/Deep-HP (mainly written in Fortran 90) on Aurora, we have updated the NVIDIA-based GPU code from OpenACC/CUDA to OpenMP/SYCL. This direct strategy preserves the initial code structure, paths the way to portability across multi-platform, minimizes re-implementation effort, and technically keeps intact the global optimization. On one hand, the translation from OpenACC/CUDA to OpenMP/SYCL implementation preserves control over data transfers between CPUs and GPUs with minor exceptions, as well as parallelization of the computationally intensive kernels ; while on the other, distributed parallelism through the native 3D domain decomposition of Tinker-HP[50, 1] can be directly exploited through MPI calls between Intel Xe GPUs. Contrary to our simulations on Nvidia GPUs, we resorted to an Aurora-compatible version of Pytorch[66] and not JAX for the computation of the neural network potential energy and forces. Following the same protocol described on Jean Zay (see previous section), we obtain 0.03 ns/day (mixed precision) per single adaptive sampling replica (32 nodes, 384 GPUs).As expected since this is a new portage of the code that will benefit of further optimization over the next months, the present production per replica on Aurora is lower than the numbers per replica obtained on Jean Zay which used the reference Tinker-HP implementation. One constraint is related to memory: Jean Zay’s H100 propose 80Gb/GPU while Aurora is adding 64Gb per Xe GPU. However, Aurora benefits from its Exascale capabilities and from its associated extremely large availability of computing nodes. Aurora can achieve a similar performance as Jean Zay with 15 replica (5760 GPUs, 480 nodes).
@@
Using the STMV simulations, we open a new window on the impact of pH fluctuations on biological systems, particularly viruses, which is a critical factor influencing their structural dynamics and function. Traditional molecular dynamics (MD) approaches often simulate pH effects by manually adjusting the protonation states of ionizable residues to match the desired pH conditions or rely on specific constant-pH MD implementations[17] that are strongly dependent on the force field quality. For STMV, as for any other viral particle, pH-dependent changes play a key role in the infection process, likely affecting capsid stability and RNA-capsid interactions. Simulating the pH transition from physiological to more acidic reflects conditions likely to be encountered during the early stages of viral cell entry and intracellular transport[84]. Implying the ”beyond DFT” Foundation Neural Network, we simulated the transition from pH$=$7 to pH$=$5 by explicitly introducing the calculated number of free protons into the system, based on pKa estimations. This approach allows proton exchange to occur dynamically and interactively with the molecular environment. As a proof of concept, we extracted a viral STMV capsomer (33,317 atoms) and quantitatively estimated the number of proton transfer events related to the 6-ns capsomer simulation (Fig. 4), providing a dynamic portrait of the molecular response to acidification.
@@
VI Conclusion
We presented an end-to-end integrated computational strategy for generating synthetic quantum chemistry datasets that significantly enhance the accuracy and predictive power of atomistic Foundation Machine Learning models. Our method pushes the boundaries of quantum chemical calculations, especially Diffusion Quantum Monte Carlo and selected Configuration Interaction, to unprecedented scales. Leveraging the power of exascale computing resources, we computed energies and forces with near-exact quantum accuracy at the complete basis set limit for tens of thousands of molecular configurations, establishing a new state-of-the-art in dataset generation. In a near future, going beyond DMC and the combination of multi-determinant QMC energies and forces with selected-CI wavefunction could be difficult with classical computers. However, one next step for constructing such high accuracy dataset could be to complement it with exact Full-CI/full basis quality quantum computing results (see Figure 1) once logical qubits become available. As an Easter egg, we already included in the dataset a dozen ”chemically accurate” converged energy datapoints (i.e. equivalent to a (sCI+PT2),full basis result) obtained with our GPU-accelerated quantum emulator and extracted from a recent study[95]. This should generalize in the future.
Key algorithmic and computational innovations include a highly optimized multi-level batching scheme, a novel unified memory management strategy, and the optimization for GPUs of zero-variance force estimators within the QMC method. Together, these advancements enabled efficient and scalable calculations on leadership-class computing facilities such as Aurora. Specifically, our optimized GPU implementations achieved unparalleled performance, reaching sustained multi-petaflops computational throughput and overcoming significant bottlenecks traditionally associated with QMC force computations.
We demonstrated the practical applicability of our computational pipeline using FeNNix-Bio1, a new generation foundation neural network model trained with our high-fidelity datasets including DMC results. Coupled with path integral-based Ring Polymer Molecular Dynamics (RPMD) and adaptive sampling, FeNNix-Bio1 successfully simulated condensed-phase quantum dynamics of a complex biological system comprising one million atoms with 32 RPMD beads to fully include nuclear quantum effects. This capability provides an unprecedented view into biochemical processes involving bond formation, proton transfers, and chemical reactions in realistic biosystems at quantum chemical accuracy. It is important to note that thanks to the generated dataset, the foundation model is not the only product that can be directly generated. Indeed, thanks to a recently proposed directed graph attention networks approach[19], the discussed FeNNol library can generate, within hours, new version of classical force fields such as GAFF (Generalized Amber Force Field)[101] that can be used to prepare initial simulation setups at very limited computational cost. This work signifies a major step forward, bridging the gap between chemically accurate quantum methods and molecular dynamics simulations for biologically relevant scales. The ability to simulate multi-nanosecond quantum molecular dynamics on such large systems including their chemical reactivity thanks to a Foundation ML Model will profoundly impact our understanding of fundamental biochemical processes, potentially revolutionizing fields such as drug discovery, enzyme catalysis, and biomolecular engineering. Future work will explore even larger scales and more diverse chemical spaces, laying the foundation for general-purpose ML potentials of unprecedented accuracy and versatility towards predictive molecular modeling in Chemistry and Biology.

### v3

@@ V-B1 QMC Energy and Wavefunction Evaluation
We developed a highly optimized implementation of QMC calculations in QMCPACK that leverages the massive parallelism of exascale hardware. Building on previous optimizations for material science applications, we specifically enhanced the Linear Combination of Atomic Orbitals (LCAO) branch to accelerate molecular applications on GPU architectures. Our implementation enables unprecedented calculation throughput, processing over 40,000 molecular configurations with QMC (20,251 with forces) on the Aurora supercomputer, achieving 124 PFLOPS sustained performance on 2048 nodes, and while using over 4500 walkers per rank (see Fig4).
@@
These implementations reduced kernel launch overhead by an order of magnitude and significantly improved memory throughput, which is particularly important for the LCAO approach used in molecular systems where the number of electrons is small. The GPUs’ architecture-independent optimizations achieved 60% of theoretical peak performance for typical molecular configurations in our dataset.
@@ V-B2 Single Determinant QMC Forces
For the QMC layer of our quantum chemistry database generation protocol, we build upon the existing force evaluation formalism in QMCPACK, developed according to Filippi et al.’s  [29] approach. Rather than re-implementing the mathematical framework, we focused on optimizing the computational efficiency for exascale architectures by integrating the force evaluation with QMCPACK’s batched structure. This optimization enabled us to compute accurate interatomic forces for over 20,000 molecular configurations, achieving 62 PFLOPS sustained performance on Aurora with scaling efficiency of 30% up to 2048 nodes.
@@
By integrating force calculations with the batched structure of QMCPACK, we eliminated redundant computations and maximized hardware utilization, enabling force calculations at scales previously impossible with QMC methods. The performance improvements from our batched implementation make QMC force calculations practical for large-scale applications, opening new possibilities for highly accurate MD simulations and foundation model training based on QMC-level accuracy.
@@ V-B3 Multideterminant QMC Forces
  • Utilizing the same flattened center indexing scheme used elsewhere in the code.
@@
  • Minimize data transfers between CPU and GPU;
  • Ensure optimal data locality for GPU processing;
  • Reuse memory allocations across different parts of the calculation.
@@
This implementation makes multideterminant QMC forces practically usable at scale for the first time. By completing, optimizing, and integrating this capability with our batched computation structure, we have created a pathway for generating higher-accuracy reference data for ML potentials, particularly for systems where electronic correlation effects are significant.
@@ V-B5 QMC Submission Engine on Aurora
QMCPACK calculations on Aurora were facilitated by a custom two-level job submission engine built on top of the Nexus[100] workflow system. To enable to maximize computational efficiency, molecular systems were batched together by electron count similarity as this supports greater execution uniformity among batched molecules. This top level parallelism, across batches, was complemented by further separation of molecules into sub-ensembles, each of which ran under a single top-level QMCPACK MPI process to maximize robustness of the simulations, providing a simple form of fault tolerance. The balance between these levels, as well as assignment of individual molecules to selected node resources, was optimized to simultaneously maximize job submission size, fault tolerance, and computational efficiency.
VI Performance Measurement
To demonstrate the computational efficiency and accuracy of our new QMC implementation, we performed extensive performance analyses on several representative molecular systems. We primarily focused on water molecules (monomer, dimer, and trimer) due to their biological relevance and the well-known computational challenges they pose for accurate quantum chemical treatment. Additionally, we benchmarked a representative SPICE molecule with heavier atoms containing 16 electrons for force evaluation, as well as a larger, computationally demanding molecule with 176 electrons to examine scalability.
Figure 2 presents a comprehensive comparison between the DFT forces and the QMC mixed estimator forces for a representative dataset. In these plots, the x-axis represents the DFT force components (in Ha/bohr), while the y-axis corresponds to the mixed estimator forces, also in Ha/bohr which is computed via the mixed estimator defined as $F{mixed}=2F{DMC}-F{VMC}$. Across the three Cartesian components (x, y, z), a strong linear correlation is observed between the DFT forces and the QMC mixed estimator $F_{mixed}$ as demonstrated by the clustering of points along the ideal diagonal line (shown as a dashed line in each subplot). The slight systematic differences are present between the DFT values and the Mixed estimator forces show the subtle differences between the methods and reflecting the increase of accuracy.
Figure 2: Comparison of force components between DFT and QMC mixed estimator. The three different marker shapes and colors represent the Cartesian components of the forces. The forces correlate with the DFT very well while the DFT shows significant discrepancies due to the lack of accuracy of the method.
VII Performance Results
VII-1 Single-Determinant QMC Performance
@@
Figure 4 illustrates the substantial performance gains achieved with our new algorithm. We observe clear improvements across water molecules and clusters with electron counts ranging from 8 to 152 electrons. Quantitative speedups measured against the CPU baseline are summarized in Table I, highlighting consistent and significant GPU acceleration.
@@
VII-2 Resource Allocation Strategy
@@ VII-3 Multideterminant QMC Performance
We successfully executed multideterminant QMC calculations for the water monomer; however, performing analogous calculations on the water dimer exceeded our computational allocation limits, underscoring the extreme intensity of such computations. To achieve optimal performance with multideterminant wavefunctions, we allocated approximately 20 Aurora nodes per molecule and processed batches of 100 molecules per job (utilizing 2000 nodes per job submission). Our implementation reached a sustained computational throughput of 24 TFLOPS per node configuration on Aurora, corresponding to 25% of the peak theoretical performance. These substantial accuracy gains justify the computational costs incurred, demonstrating the critical importance of multideterminant QMC methods in achieving chemically accurate reference datasets for training next-generation ML models.
VIII Implications
We presented an end-to-end integrated computational strategy for generating synthetic quantum chemistry datasets that significantly enhance the accuracy and predictive power of atomistic Foundation ML models. Our method pushes the boundaries of quantum chemical calculations, especially Diffusion QMC and selected Configuration Interaction, to unprecedented scales. Leveraging the power of exascale computing resources, we computed energies and forces with near-exact quantum accuracy at the CBS limit for tens of thousands of molecular configurations, establishing a new state-of-the-art in dataset generation.
Key algorithmic and computational innovations include a highly optimized multi-level batching scheme, a novel unified memory management strategy, and the optimization for GPUs of zero-variance force estimators within the QMC method. Together, these advancements enabled efficient and scalable calculations on leadership-class computing facilities such as Aurora. Specifically, our optimized GPU implementations achieved unparalleled performance, reaching sustained multi-petaflops computational throughput and overcoming significant bottlenecks traditionally associated with QMC force computations. Improving further the accuracy beyond the present DMC and the combination of multi-determinant QMC energies and forces with selected-CI wavefunctions could be difficult with classical computers due to the previously discussed Exponential Wall problem. However, once logical qubits become available[101], Quantum Computing could be a solution to overcome this barrier and bring, at least, useful polynomial speedups[102]. As this would allow access to exact Schrödinger-like Full-CI/CBS computations (Figure 1), we already considered these new class of quantum algorithm for chemistry in the construction of our present dataset. As an Easter egg, we included a dozen datapoints, converged at the CBS chemical accuracy that were obtained with our GPU-accelerated quantum emulator[103]. This should generalize in the future as the algorithms, quantum emulators and quantum hardware mature. Overall, we demonstrated the practical applicability of our computational pipeline using the FeNNix-Bio1 foundation NNP model trained with our high-fidelity datasets. Therefore large scale applications became possible.
VIII-A Large Scale Quantum Dynamics Simulations with a ”beyond DFT” Foundation Neural Network Model
@@
Using the CBS QMC and sCI data produced in the project, we improved an early version of the FeNNix-Bio1 model [81] using transfer ML via the FeNNol library[78] to obtain a ”beyond DFT” improved approach. We then coupled the improved model to the GPU-accelerated massively parallel Tinker-HP MD package[104, 105] and leverage simultaneously its Deep-HP (neural networks), Quantum-HP (quantum nuclear effects) modules while using enhanced sampling via non-supervised adaptive sampling MD. We tested the new approach on both Jean-Zay (H100) and Aurora(Xe) machines on the full structure of the Satellite Tobacco Virus, a 1-million atom virus. The simulation encompassed 32-million particles (1-million atom X 32 beads) since it leverages Ring Polymer quantum MD simulation. We extended the conformational space sampling with adaptive sampling replica to decompose the global exploration problem into a set of separate MD trajectories. These latter can be restarted thanks to a selective process in order to achieve sufficient phase-space sampling.
@@
As a proof of concept, in order to improve further the quality of the DFT-based FeNNix-Bio1 model towards a ”beyond DFT” model, we leveraged transfer learning,[106] to include the high-quality, CBS, DMC energies and forces present in the SPICE2(+)-ccECP dataset. In practice, starting from the model trained on the DFT part of the dataset, we reoptimize the energy heads (freezing the parameters of the embedding), on the QMC energies and forces with the same loss function as in the original training. As is common practice, we started with a smaller learning rate of $10^{-5}$ using the adabelief optimizer[107]. Note that the QMC energies and forces inherently contain statistical noises, with a significantly larger variance for the forces. We take this into account in the loss function that we optimize for transfer learning by putting 100 times more weight on the energies than on the forces. Starting from a MAE of 15.3 kcal/mol (RMSE of 16.3 kcal/mol) on the energies and MAE of 5.7 kcal/mol/Å (RMSE of 8.7 kcal/mol/Å) on the forces, we end up, on the validation set, after 1000 epochs of training, with a MAE of 0.54 kcal/mol (RMSE of 2.4 kcal/mol) on the energies and MAE of 4.6 kcal/mol/Å (RMSE of 6.3 kcal/mol/Å) on the forces. These results are shown in Figure 5.
Figure 5: Distribution of the energy and force components errors on the QMC validation set, before and after transfer learning
We assess the changes in the model due to the transfer learning by computing the interaction energies of low-lying water clusters up to hexamers[108, 109]. Compared to CCSD(T)/CBS energies, we observe a systematic improvement of the machine learning model as shown in Figure 6. Starting from a MAE of 0.38 kcal/mol (RMSE of 0.5 kcal/mol), we end up with a MAE of 0.20 kcal/mol (RMSE of 0.24 kcal/mol).
Beside accumulating more QMC data to better cover chemical space, future work will explore how to exploit the variance estimation provided by QMC in the transfer learning loss function.
Figure 6: Per molecule interaction energy on the low-lying water clusters, before and after transfer learning, compared to CCSD(T)/CBS[108, 109]
As a note, it is important to quote that thanks to the generated dataset, the foundation model is not the only product that can be directly generated. Indeed, thanks to directed graph attention networks[110], the discussed FeNNol library can generate, within hours, new version of classical force fields such as GAFF (Generalized Amber Force Field)[111] that can be used to prepare initial simulation setups at very limited computational cost.
@@
We coupled our newly developed JAX-based [112] FeNNoL library [78] that allows for the construction and inference of atomistic ML potentials such as FeNNix-Bio1[81] to the Tinker-HP MD package[104, 105] via the Deep-HP interface[74]. By doing so, we leverage the Ring Polymer MD (RPMD)[113] capabilities of Tinker-HP (Quantum-HP module [114]) in combination with FeNNix-Bio1 in order to recover quantum dynamics and statistics of the Satellite Tobacco Virus (STMV)[115]. STMV is a good model system for studying viral assembly and structural dynamics as it is includes a single-stranded RNA genome located within an icosahedral capsid. RPMD can be used efficiently for determining accurate properties of condensed-phase molecular systems[2, 8]. It uses classical statistical physics to mimic quantum dynamics via fictitious ring polymer which consist in n copies (or beads) of the system particles connected by harmonic springs[114]. Obtaining large scale reference simulation data using RPMD on meaningful biological systems will help to improve the present FeNNix-Bio1 near-classical computational cost quantum thermostats such as the Adaptative Quantum Thermal Bath[8] that can already generate orders of magnitude longer simulations but that are more approximate. To ensure full accuracy at room temperature, we used 32 beads per atom in RPMD simulations, a number where simulations are essentially converged[2, 8]. These latter leverage two levels of parallelism: one at the beads level and a second one within each bead. Additionally, to push further the sampling of the conformational space, we added on top of the RPMD framework an additional layer of parallelism using unsupervised adaptive sampling[116]. Here multiple iterations of independent MD simulations replica are launched to ”mine” the conformational space. The selection of the initial structures at each iteration follows an adaptive, fully automated, procedure enhancing the exploration of a low-dimensional space of slow variables and maximizing the conformational space exploration by penalizing areas that have already been extensively visited. This unsupervised selection step has the advantage of suppressing the problem of the choice of the initial collective variable at the beginning of the simulation. In practice the number of replica is only limited by the number of available GPUs, making adaptive sampling particularly suited for exascale computing[116]. On Jean-Zay, each of the bead was parallelized on 8 H100 GPUs for a total of 8x32=256 H100s. Exploiting parallelism within each bead is mandatory to fit in the memory of each GPU thanks to memory distribution. All in all, we obtain 0.11 ns/day (mixed precision) per single adaptive sampling replica, making production of tens of nanoseconds per day of such quantum dynamics simulation accessible through adaptive sampling given enough resources. To our knowledge, achieving such level of quantum dynamics production with a foundation model grounded on extremely accurate quantum chemistry dataset on such a large biosystem is unprecedented. The Jean Zay machine having 1280 GPUs, we were able to launch 4 simultaneous replica leveraging 1024 GPUs, producing therefore 0.44 ns/day of adaptive quantum dynamics. Note that such adaptive setup can also be used sequentially to extend the number of replica that can be launched one after another enabling to accumulate simulation time. Finally, further H100 optimization of the combined computing layers should possibly enable us to achieve the 1 ns/day threshold with the same number of replica whereas adding more replica is also a solution. This motivated our portage on the Exascale-class Aurora machine.
@@
In the perspective to run Tinker-HP/Deep-HP (mainly written in Fortran 90) on Aurora, we have updated the NVIDIA-based GPU code from OpenACC/CUDA to OpenMP/SYCL. This direct strategy preserves the initial code structure, paths the way to portability across multi-platform, minimizes re-implementation effort, and technically keeps intact the global optimization. On one hand, the translation from OpenACC/CUDA to OpenMP/SYCL implementation preserves control over data transfers between CPUs and GPUs with minor exceptions, as well as parallelization of the computationally intensive kernels ; while on the other, distributed parallelism through the native 3D domain decomposition of Tinker-HP[104, 105] can be directly exploited through MPI calls between Intel Xe GPUs. Contrary to our simulations on Nvidia GPUs, we resorted to an Aurora-compatible version of Pytorch[117] and not JAX for the computation of the NNP energy and forces. Following the same protocol described on Jean Zay (see previous section), we obtain 0.03 ns/day (mixed precision) per single adaptive sampling replica (32 nodes, 384 GPUs).As expected since this is a new portage of the code that will benefit of further optimization over the next months, the present production per replica on Aurora is lower than the numbers per replica obtained on Jean Zay which used the reference Tinker-HP implementation. One constraint is related to memory: Jean Zay’s H100 propose 80Gb/GPU while Aurora is adding 64Gb per Xe GPU. However, Aurora benefits from its Exascale capabilities and from its associated extremely large availability of computing nodes. Aurora can achieve a similar performance as Jean Zay with 15 replica (5760 GPUs, 480 nodes).
@@
The simulation protocol was chosen similar as in ref [81] (see section V. B). Using the STMV simulations, we open a new window on the impact of pH fluctuations on biological systems, particularly viruses, which is a critical factor influencing their structural dynamics and function. Traditional MD approaches often simulate pH effects by manually adjusting the protonation states of ionizable residues to match the desired pH conditions or rely on specific constant-pH MD implementations[118] that are strongly dependent on the force field quality. For STMV, as for any other viral particle, pH-dependent changes play a key role in the infection process, likely affecting capsid stability and RNA-capsid interactions. Simulating the pH transition from physiological to more acidic reflects conditions likely to be encountered during the early stages of viral cell entry and intracellular transport[119]. Implying the ”beyond DFT” Foundation Neural Network, we simulated the transition from pH$=$7 to pH$=$5 by explicitly introducing the calculated number of free protons into the system, based on pKa estimations. This approach allows proton exchange to occur dynamically and interactively with the molecular environment. As a proof of concept, we extracted a viral STMV capsomer (33,317 atoms) and quantitatively estimated the number of proton transfer events related to the 6-ns capsomer simulation (Fig. 7), providing a dynamic portrait of the molecular response to acidification.
@@
Perspectives We demonstrated the practical applicability of our computational pipeline using FeNNix-Bio1, a new generation foundation NNP model trained with our high-fidelity datasets including QMC/s-CI results. Coupled with RPMD and adaptive sampling, the present implementation will be further optimized allowing to perform million-atom biosystems condensed-phase quantum dynamics simulations approaching the full quantum chemical accuracy. This will enable a closer look at biochemical processes involving chemical reactions. This major step forward bridges the gap between chemically accurate quantum methods and ML simulations for biologically relevant scales. The ML ability to simulate multi-nanosecond quantum MD on such large systems including their chemical reactivity will profoundly impact our understanding of fundamental biochemical processes, potentially revolutionizing fields such as drug discovery, enzyme catalysis, and biomolecular engineering. Future work will explore even larger scales and more diverse chemical spaces, laying the foundation for improved general-purpose ML potentials of unprecedented accuracy and versatility towards predictive molecular modeling in Chemistry, Biology and Drug Design.